OpenStudy (anonymous):

What is the simplified form of 9 times y to the seventh power over 15 times x to the fifth power divided by 5 x squared over 45 y cubed ?

OpenStudy (saifoo.khan):

\[\frac{9y^7}{15} \times x^5 \div \frac{x^2}{45}\]Like that?

OpenStudy (anonymous):

9y^7 45y^3 27y^10 ------ * --------- = --------- 15 x^5 5x^2 5x^7